10 ml of N-HCL, 20 ml of N/2 H2SO4 and 30 ml of N/3 HNO3 are mixed together and volume made to one litre. The normality of H+ in the resulting solution

Grade:12th pass

3 Answers

Khushi
5 years ago
Meq=N×V 
So for HCl=1×5=5M
For H2SO4=1/2 ×20=10M
For HNO3=1/3×30=10M
Mtotal=M1+M2+M3= 5+10+10=25M
volume =1000mL
Meq=N×V
25=N×1000
25N/1000
=N/40
Raj Ranjan Dey
5 years ago
Eq= N x V 
For HCl
Eq= 1 x 10 = 10 Eq
 
For H2SO4
Eq = 1/2 x 20 = 10 Eq
 
For HNO3
Eq= 1/3 x 30 = 10 Eq
 
Total Eq = 10+10+10 = 30 Eq
 
Again, 
Eq= N x V 
30 = N x 1000 
3N/100 Ans
Nisha Sangwan
2 years ago
H2SO4=1/2 ×20=10M
HNO3=1/3×30=10M
Mtotal=M1+M2+M3= 5+10+10=25M
volume =1000mL
Meq=N×V
25=N×1000
25N/1000
=N/40